How many g of gold should a coin of 59% gold be it when combined with a 9 g pure gold necklace it forms a metal that is 74% gold?
Answer by jorel1380(3719)

Let g be the weight of the coin, which is 59% gold. Then:
.59g + 9=0.74(g+9)
.59g+9=0.74g+6.66
.15g=2.34
g=15.6 gm as the weight of the coin
☺☺☺☺